OpenStudy (anonymous):

"Three people who work full-time are to work together on a project, but their total time on the project is to be equivalent to that of only one person working full-time. If one of the people is budgeted for one-half of his time and a second is budgeted for one-third her time, what portion of the third worker's time should be budgeted to this project?": I've gotten as far as 1/2+1/3+?=1whole...how would I calculate it without a calculator?

jimthompson5910 (jim_thompson5910):

so it looks like you want to solve 1/2 + 1/3 + x = 1 for x

OpenStudy (anonymous):

right

jimthompson5910 (jim_thompson5910):

1/2 + 1/3 + x = 1 3/6 + 2/6 + x = 1 (3 + 2)/6 + x = 1 5/6 + x = 1 x = 1 - 5/6 x = 6/6 - 5/6 x = (6 - 5)/6 x = 1/6

jimthompson5910 (jim_thompson5910):

So the third worker should get 1/6 of the job

OpenStudy (whpalmer4):

\[\frac{1}{2} + \frac{1}{3}+x = 1\]LCM of 2 and 3 is 6, so multiply everything by 6. \[3+2+6x=6\]\[6x=1\]\[x=\frac{1}{6}\]

OpenStudy (whpalmer4):

There's an assumption made here in the problem which doesn't really hold up well in real life, namely that everyone works the same amount of time in a "full day"...

OpenStudy (anonymous):

Thank you both so much!! best answers I've had on any questions!! I really appreciate it!

OpenStudy (whpalmer4):

You're welcome!

jimthompson5910 (jim_thompson5910):

glad to be of help

OpenStudy (anonymous):

I know its an inconsistent question...I got it from a samle Accuplacer test to study... :/

OpenStudy (whpalmer4):

Someone looking for extra work to get better? Wow, quite a refreshing change from all too many OpenStudiers who just want answers to finish their test or homework!

OpenStudy (anonymous):

HAHA nah, I'm studying to take a placement test...I have the answers (to be fair, so I know if some one shoots me a b.s. answer... thanks some people) I just want to know how to get the job done for myself. Which is why I really appreciate your explanation!

jimthompson5910 (jim_thompson5910):

that's good to know since you at least have some idea of where you're going if anyone just gave you the answer, just report them because it's against the code of conduct anyway (you have to show how you got it or help the other person get it). It doesn't help anyway with just the answer.

OpenStudy (anonymous):

Exactly! And when I asked them to explain, they didn't. So my first question was never really answered...but I took it down cause I thought 'what's the point'? I don't want to sit all night waiting for ONE answer.

jimthompson5910 (jim_thompson5910):

hmm well some unfortunately can be rude like that, so it's best to move on

OpenStudy (anonymous):

Right, thanks again!

jimthompson5910 (jim_thompson5910):

you're welcome